OpenStudy (anonymous):

9x^2+x=5 this is the quadtratic formula: can someone help get is set up to solve

OpenStudy (anonymous):

9x^2+x-5=0

OpenStudy (anonymous):

and use the quadratic equation

OpenStudy (anonymous):

ax^2 + bx + c = 0

OpenStudy (anonymous):

so the a= b=1 c=5?

OpenStudy (anonymous):

a=9?

OpenStudy (anonymous):

c=-5

OpenStudy (anonymous):

x=( -b +or - sqrt( b^2 - 4 * a *c))/2*a

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But c=-5

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Since quadratic equation should be of the form ax^2+bx+c

OpenStudy (anonymous):

the plus or minus means you have to do the equation twice one with + and one with - which means you should have two answers for x

OpenStudy (anonymous):

ax^2 + bx + c = 0 so \[x=\frac{-b \pm \sqrt{b ^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

amir looks better follow that one

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

i was makin that equation and it take a while if it didn't i'll answer first

OpenStudy (anonymous):

amir.sat will you verify my answer is correct once i work the problem?

OpenStudy (anonymous):

where's you answer?

OpenStudy (anonymous):

\[-1\pm \sqrt{9 /18}\]

OpenStudy (anonymous):

or does the 9/18 reduce futher

OpenStudy (anonymous):

what's this? is it the solve way of that problem?

OpenStudy (anonymous):

replace the value in the formula that i've posted.

OpenStudy (anonymous):

a=9 b=1 c=-5? right I just plug in the numbers

OpenStudy (anonymous):

so you are able to find the values of x

OpenStudy (anonymous):

what is your answer? please

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

\[x=\frac{-1+\sqrt{1+1800}}{18}\] \[x=\frac{-1-\sqrt{1+1800}}{18}\]

OpenStudy (anonymous):

-4ac= -4(9)(-5) = 180?

OpenStudy (anonymous):

not sure where you got 1800

OpenStudy (anonymous):

yes okk it's 180

OpenStudy (anonymous):

and the square root of 180 =\[6\sqrt{5}\]

OpenStudy (anonymous):

no it's 180+1

OpenStudy (anonymous):

okay then it's 9/18

OpenStudy (anonymous):

it's not